horizontal escalator connecting airport terminals is 1km long. two kids from opposite ends and run toward each other at 3m/s and meet 600m from the start of the escalator. what is the speed of the escalator?
1
Expert's answer
2020-04-02T10:43:37-0400

So,


"(v+v_0)t=600" "\to" "t=\frac{600}{v+v_0}"


"(v-v_0)t=400" "\to" "t=\frac{400}{v-v_0}"


"\frac{600}{v+v_0}=\frac{400}{v-v_0}\to\frac{600}{3+v_0}=\frac{400}{3-v_0}"


"v_0=\frac{600}{1000}=0.6" "m\/s" Answer